OpenStudy (anonymous):

find cost of painting the sides of a water tank measuring 225cm in diameter and 375cm in height at P20.00 per square meter.

Directrix (directrix):

Lateral Area = 2*pi*r*h where r is the radius of the circular base and height is the height of the cylinder. LA = 2 pi r h LA = 2 pi (225/2)(375) LA = 225 pi (375) LA = 265071.880147 square cm = 26.507188m² Cost at $20 per square meter is 26.507188 *20 = $ 530.14